Angry​ Birds, Inc. designs and builds retaining walls for individual customers. On August​ 1, 2 jobs were in​ process: Level 1 with a beginning balance of​ $12,000 and Level 2 with a beginning balance of​ $8,600. During​ August, Level 3 and Level 4 were started. Levels 1 and 2 were sold at cost plus 30​% during August. All other jobs remained in process at month end. The company applies manufacturing overhead at a rate of ​$9 per direct labor hour. Direct labor wages average ​$10 per hour. The following additional cost information is available for​ August:

Level 1

Level 2

Level 3

Level 4

Direct Materials

​$1,200

​$8,000

​$2,100

​$3,000

Direct Labor

​$1,800

​$4,800

​$360

​$1,200

The sales staff is paid a​ 2% commission on all sales for the period. The company closes out the total​ over/underapplied manufacturing overhead to Cost of Goods Sold at the end of each month. Calculate the following given actual MOH of​ $8,200 incurred during August. The amount of MOH applied to all jobs in process during August equals ​$ nothing. The adjusted COGS reported on the Income Statement for August equals ​$ nothing. The gross profit reported on the Income Statement for August equals ​$ nothing. The ending work in process balance reported on the Balance Sheet for August equals ​$ nothing. Round final answers to the nearest dollar.
Answer #1
The amount of MOH applied to all jobs in process during August :
Level 1 Level 2 Level 3 Level 4 Total
Direct labor a 1800 4800 360 1200
Direct labor rate per hour b 10 10 10 10
Direct labor hours c=(a/b) 180 480 36 120
Predetermined overhead rate d 9 9 9 9
MOH applied (c*d) 1620 4320 324 1080 7344
The adjusted COGS reported on the Income Statement for August:
Adjusted cost of goods sold:
$
Cost of goods sold (Note:1) 42340
Add: Under absorbed overhead 856
Adjusted cost of goods sold 43196
Note:1- cost of goods sold
Cost of goods sold=Cost of jobs sold=Cost of level 1 and 2
Level 1 Level 2 Total
Beg. Bal 12000 8600
Direct Materials 1200 8000
Direct labor 1800 4800
MOH applied (Refer prev. part) 1620 4320
Total of above=Cost of goods sold 16620 25720 42340
Total MOH applied to jobs=$ 7344
Actual MOH=$ 8200
Applied MOH < Actual MOH, Hence under-absorption
Under absorbed overhead=Actual MOH-Applied MOH=8200-7344=$ 856
The gross profit reported on the Income Statement for August:
$
Sales revenue (Note:2) 55042
Less:Adjusted cost of goods sold 43196
Gross profit 11846
Note:2 - Sales revenue
Sales revenue=Cost of jobs sold+30%=42340+30%=$ 55042
The ending work in process balance reported on the Balance Sheet for August:
Ending work in process balance=Cost of jobs incomplete=Cost of level 3 &4
Level 3 Level 4 Total
Direct Materials 2100 3000
Direct labor 360 1200
MOH applied (Refer first part) 324 1080
Total of above=Ending WIP 2784 5280 8064